OpenStudy (anonymous):

Help me please! Solve? -4(n – 6) = 12 A.) n=-4.5 B.) n=3 C.) n=9 D.)n=22

OpenStudy (anonymous):

The answer is B. let me show you how

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yes!(:

OpenStudy (anonymous):

First, factor out -4(n-6). Which is -4*n= 4n. And -4*-6= 24

OpenStudy (anonymous):

Then subtract the 24 to the other side of the equation. Because we want like terms together. -4n= 12-24. That becomes -4n=-12 Divide both sides by -4 -4 n/-4= n. then -12/-4= 3 So n= 3

OpenStudy (anonymous):

okay thanks(:

OpenStudy (anonymous):

Do you understand it know? Just make sure you pay attention to the signs ok. :)

OpenStudy (anonymous):

yes i do(: